OpenStudy (math2400):

QUANTITATIVE If one species (2n = 16) crosses with another species (2n = 22), producing an allopolyploid offspring, what is the ploidy of the offspring?

OpenStudy (anonymous):

2 different species with different number of chr cannot cross

OpenStudy (math2400):

that's what i thought but we are only given the answer 2n = 8 + 11 = 19 2n = 22 4n = 32 + 44 = 76 2n = 16 and I'm so sorry i just realized i posted in math>< i was in biology the last time i was logged in